# Prediction Sets and Conformal Inference with

### v2

@@ 2 Setup and definitions
The oracle prediction set under point identification is $C^{*}_{P_{Y|X}}=\left\{(x,y):x\in\mathcal{X},y\in L(x,\lambda_{x}^{\alpha})\right\}$ which is both conditionally and marginally valid. For a given $x$, it collects the values of $y$ with sufficiently high density and intuitively assembles the set of $y$’s with the highest probability for a given volume. Once this is calibrated to yield the desired miscoverage level, one obtains a minimal-volume set. Estimating the oracle prediction set $C^{*}_{P_{Y|X}}$ therefore reduces to the problem of estimating the upper level sets $L(x,\lambda)$ of the conditional density $p(y\mid x)$ and selecting an appropriate threshold parameter.
@@
Suppose $Y$ is interval censored, for any $\lambda<\infty$ and any $y\in[a,b]$, where $a<b$ and $(a,b)\in\mathbb{R}^{2}$ is in the support of $(Y^{L},Y^{U})$, there exists some $P^{\prime}\in\mathcal{P}_{I}$, such that $y\in L^{\prime}(\lambda)$, where $L^{\prime}(\lambda)$ is the upper level set for the density $p^{\prime}(y)$ of $P^{\prime}$.
Proposition 2.1 states that for any point $y\in[a,b]$ such that $(Y^{L},Y^{U})$ has positive density at $(a,b)\in\mathbb{R}^{2}$, we can find a random variable $Y^{\prime}$ that is observationally equivalent to $Y$ under censoring, and this $Y^{\prime}$ admits a level set that contains $y$ for any threshold $\lambda$. This is possible because one can adjust the density of $Y^{\prime}$ at any point $y$ satisfying the stated conditions to be arbitrarily large while ensuring that this density remains within the identified set $P_{I}$. [^3] For example, consider a random variable $Y$ with support on $\mathbb{R}$ that is censored within $[0,1]$: we do not observe the value of $Y$ when it lies in $[0,1]$ but only observe the indicator $1[Y\in[0,1]]$. In this case, we cannot reject the hypothesis that any value $y\in[0,1]$ belongs to the level set of $Y$. This lack of identification for level sets is due to the possibility of pointwise pathological behaviour of the density function.
In contrast to the inability to bound the density function, the following two lemmas characterise the identified set of the conditional distribution of $Y$ given $X$ under interval censoring. Lemma 2.2 relates the sharp identified set of the conditional distribution of $Y$ given $X$ to the conditional distribution of $(Y^{L},Y^{U})$, which can be directly identified and estimated from the data, and provides a feasible lower bound for the restriction in Equation (4). Lemma 2.3 shows that the lower bound is tight, and we are not unnecessarily conservative in constructing the oracle prediction set.
Lemma 2.2 (Theorem SIR-2.3 in molinari2020ChapterMicroeconometrics).
@@
Lemma 2.2 suggests that we could implement a feasible version of Equation (4), by replacing the optimisation constraint $\inf_{P\in\mathcal{P}_{I}}P\left(Y\in C(x)\mid X=x\right)\geq 1-\alpha$ with $P\left([Y^{L},Y^{U}]\subset C(x)\mid X=x\right)\geq 1-\alpha$, which implies the original constraint, since for $C\in\mathcal{C}$, we can write $C=\sqcup_{m=1}^{M}[a_{m},b_{m}]$
@@ 3 Consistency
We first consider the estimation of the oracle prediction interval $C^{*}_{I}=[\tau_{0}(x),\tau_{1}(x)]$, for $x\in\mathcal{X}$. This analysis highlights the conditions required to ensure the consistency of the prediction set under different schemes of interval censoring. Subsequently, we will extend our analysis to the estimation of oracle prediction sets $C^{*}\in\mathcal{C}$ consistinig of multiple intervals. To measure the difference between two sets $A,B\subset\mathbb{R}$, we use the volume of their symmetric difference, denoted by $\mu(A\bigtriangleup B)$. The symmetric difference is defined as $A\bigtriangleup B=(A\setminus B)\cup(B\setminus A)$. This is commonly used in the literature on level set estimation and conformal inference. Another commonly used notion of set distance is the Hausdorff distance. In Appendix D, we discuss the relationship between these two metrics and why the Hausdorff distance is unsuitable in certain cases.
@@
$\min_{t_{0}\leq t_{1}}\left\lvert t_{1}-t_{0}\right\rvert\quad\text{s.t.}\quad P_{n}(t_{0},t_{1};x)\geq 1-\alpha.$ (7)
@@
where $K_{h}(x)=\frac{1}{h}K\left(\frac{x}{h}\right)$ for a choice of kernel smoothing function $K(\cdot)$ and bandwidth parameter $h$, see tsybakov2009IntroductionNonparametric and hansen2008UniformConvergence. Another candidate is the distributional regression estimator [see, e.g.,][]foresi1995ConditionalDistributionb, chernozhukov2013InferenceCounterfactual,chernozhukov2021DistributionalConformal.
@@
Assumption 3.3 is a smoothness assumption on $P(t_{0},t_{1};x)$. In particular, we are assuming that by considering a slightly larger interval $[\tau_{0}-\varepsilon,\tau_{1}+\varepsilon]$ we can strictly increase our coverage probability. Assumption 3.3 also assumes an upper bound on the desity, which can be violated if the conditional distribution of $(Y^{L},Y^{U})$ is discrete, which might happen under a fixed censoring scheme.
This assumption is related to the $\gamma$-exponent condition in the level set literature, see polonik1995MeasuringMass and lei2014DistributionfreePrediction.
@@
$\min_{t_{0}<t_{1}}t_{1}-t_{0}\quad\text{s.t.}\quad P_{n}(t_{0},t_{1};x)\geq 1-\alpha-\psi_{n},$
@@
$P\left(\sup_{x\in\mathcal{X}_{n}}\sup_{t_{0}\leq t_{1}}\left\lvert P_{n}(t_{0},t_{1};x)-P(t_{0},t_{1};x)\right\rvert>\psi_{n}\right)=o(1).$
@@
Suppose $\alpha=0.5$, and $(Y_{i}^{L},Y_{i}^{U})=(0,1)$ or $(0,2)$ with equal probability, which is a discrete distribution. The oracle prediction set is $[0,1]$. With the natural estimator $P_{n}(a,b;x)=n^{-1}\sum\mathbb{I}\left\{[Y_{i}^{L},Y_{i}^{U}]\subset[a,b]\right\}$, if we solve the exact empirical analog of the oracle prediction set, as in Equation 7, we will obtain
$\check{C}=\left[0,1+\mathbb{I}\left\{\frac{1}{n}\sum Y_{i}^{U}<0.5\right\}\right],$
and the random upper bound equals $1$ or $2$ with equal probability, hence $\check{C}$ is not consistent.
@@
(1) For any $\varepsilon>0$ and $P_{X}$-almost surely $x\in\mathcal{X}$, there exists $\delta=\delta(\varepsilon)>0$, such that for any set ${C}^{\prime}(x)\in\mathcal{C}$ with $\mu({C}^{\prime}(x))\leq\mu(C^{*}(x))$ and $\mu(C^{*}(x)\triangle C^{\prime}(x))>\varepsilon$, we have $P([Y^{L},Y^{U}]\subset{C}^{\prime}(x)\mid X=x)<1-\alpha-\delta$. (2) The oracle $C^{*}(x)\in\mathcal{C}_{\bar{M}}=\{\sqcup_{m=1}^{M}[a_{m},b_{m}]:M^{\prime}\leq\bar{M}<\infty\}$ for some $\bar{M}<\infty$ and all $x\in\mathcal{X}$.
@@
4 Conformal inference with censored outcomes
@@
In particular, we will achieve finite-sample validity using the method of split conformal inference (see *romano2019ConformalizedQuantile and *lei2018DistributionfreePredictive). One feasible split conformal inference procedure for the oracle prediction interval $C^{*}_{I}$ in our setting can be described as follows. We first split the sample $\mathcal{D}$ into a training set $\mathcal{D}_{1}=\left\{Z_{i}:i\in\mathcal{I}_{1}\right\}$ and a calibration set $\mathcal{D}_{2}=\left\{Z_{i}:i\in\mathcal{I}_{2}\right\}$. We calculate the estimator $\hat{C}_{I}(x)=[\hat{\tau}_{0}(x),\hat{\tau}_{1}(x)]$ for the oracle prediction set with the training set for $x\in\mathcal{X}$ as in Equation 7. The following scores $s_{j}=s(Y^{L}_{j},Y^{U}_{j},X_{j};\hat{\tau}_{0},\hat{\tau}_{1})$ are then computed for each $j\in\mathcal{I}_{2}$ in the calibration set, given $\hat{C}_{I}(x)=[\hat{\tau}_{0}(x),\hat{\tau}_{1}(x)]$, where
@@
This score function is a modified version of the score function used in the case of conformalised quantile regression with no interval censoring studied in romano2019ConformalizedQuantile. The score function is negative when the interval $[y^{L},y^{U}]$ is contained in $[t_{0}(x),t_{1}(x)]$ and positive otherwise. By allowing for negative scores, the conformal prediction set $\tilde{C}_{I}(x)$ penalises both overcoverage and undercoverage.
@@
where $q\left(A;\zeta\right)$ denotes the $\zeta$-quantile of the set $A$ of real numbers. Since observations in $\mathcal{I}_{2}$ are assumed i.i.d., if we take another point $(X_{n+1},Y^{L}_{n+1},Y^{U}_{n+1})$, the hypothesis that it is drawn from the same distribution will not be rejected at $1-\alpha$ confidence level if $s_{n+1}\leq\vartheta_{1-\alpha}$. The conformal prediction set is then
@@
Intuitively speaking, when our estimated prediction set $\hat{C}_{I}$ is wider than necessary, the intervals $\{(Y_{j}^{L},Y_{j}^{U}),j\in\mathcal{I}_{2}\}$ tend to be subsets of $\hat{C}_{I}$ and a higher proportion of the scores computed using the calibration set will tend to be negative. In this case, the chosen quantile of the scores $\vartheta_{1-\alpha}$ is negative, and the conformal prediction interval will correct for the over-coverage of $\hat{C}_{I}$.
@@
Notice that $s_{j}\leq 0$ if and only if $[Y^{L}_{j},Y^{U}_{j}]\subset[t_{0m},t_{1m}]$ for some $m\leq M$. We can then compute $\vartheta_{1-\alpha}=q(\{s_{j}:j\in\mathcal{I}_{2}\};(1-\alpha)(1+1/\left\lvert\mathcal{I}_{2}\right\rvert))$. The conformal prediction set is then  [^4] Here, the endpoints of the estimated $\hat{C}$ are adjusted uniformly using the same $\vartheta_{1-\alpha}$. A more sophisticated adjustment method is discussed in Appendix C, where we propose a minimal Euclidean distance adjustment. We conjecture that this approach could improve the convergence rate of the conformal prediction set. The optimal choice of conformity score and conformlisation remains an interesting question.
@@
Under the assumption that $\{(X_{i},Y_{i},Y_{i}^{L},Y_{i}^{U}):i\in\mathcal{I}\cup\{n+1\}\}$ are i.i.d., the conformal prediction set $\tilde{C}$ in Equation 11 is finite-sample marginally valid, that is, for all $P$,
@@
Given a consistent estimator $\hat{C}$ for the oracle prediction set, if $s^{*}=s\left(Y^{L},Y^{U},X;C^{*}\right)$ has positive density in a neighbourhood around 00, then the conformal predictioin set $\tilde{C}$ constructed as above, for an i.i.d. sample $\mathcal{D}=\left\{\left(X_{i},Y_{i}^{L},Y^{U}_{i}\right):i\in\mathcal{I}\right\}$ which is split into training $\mathcal{I}_{1}$ and calibration sets $\mathcal{I}_{2}$, we have
@@
However, it is impossible to construct a prediction set that is finite-sample conditionally valid and consistent for the oracle prediction set lei2014DistributionfreePrediction, barber2021LimitsDistributionfree, gibbs2024ConformalPrediction, and hence we consider the weaker notion of local validity. Let $\mathcal{A}=\left\{A_{k}:k\leq K\right\}$ be a partition of $\mathcal{X}$, a prediction interval $\tilde{C}(x)$ is finite-sample locally valid with respect to $\mathcal{A}$, if for all $P$,
$\operatorname{P}\left(Y_{n+1}\in\tilde{C}_{\mathcal{A}}(X_{n+1})\mid X_{n+1}\in\mathcal{A}_{k}\right)\geq 1-\alpha.$
@@
then the local conformal prediction set is
$\tilde{C}_{\mathcal{A}}(x)=\bigcup_{m=1}^{M}\left[\hat{\tau}_{0m}(x)-\sum_{k\leq K}\mathbb{I}\left\{x\in A_{k}\right\}\vartheta_{1-\alpha,k},\hat{\tau}_{1m}(x)+\sum_{k\leq K}\mathbb{I}\left\{x\in A_{k}\right\}\vartheta_{1-\alpha,k}\right].$
The finite-sample local validity is a result of exchangeability, similar to lei2014DistributionfreePrediction.
@@
Suppose $\{(X_{i},Y_{i}^{L},Y_{i}^{U}):i\in\mathcal{I}\}$ are i.i.d. Let $\mathcal{A}$ be a partition of $\mathcal{X}$. Then the local conformal prediction set $\tilde{C}_{\mathcal{A}}(x)$ is finite-sample locally valid with respect to $\mathcal{A}$.
In the next section, we demonstrate the performance of the conformal and local conformal prediction sets in a set of numerical experiments. [^5] The Python code for implementing the proposed conformal prediction procedure, as well as replicating the numerical experiments and empirical study, is available at https://github.com/lwg342/prediction_interval_outcome.

### v3

@@ 2 Setup and definitions
The oracle prediction set under point identification is $C^{*}_{P_{Y|X}}=\left\{(x,y):x\in\mathcal{X},y\in L(x,\lambda_{x}^{\alpha})\right\}$, which is both conditionally and marginally valid (see Lei and Wasserman, (2014)). For a given $x$, it collects the values of $y$ with sufficiently high density and intuitively assembles the set of $y$’s with the highest probability for a given volume. Once this is calibrated to yield the desired miscoverage level, one obtains a minimal-volume set. Estimating the oracle prediction set $C^{*}_{P_{Y|X}}$ therefore reduces to the problem of estimating the upper level sets $L(x,\lambda)$ of the conditional density $p(y\mid x)$ and selecting an appropriate threshold parameter.
@@
Suppose $Y$ is interval valued, for any $\lambda<\infty$ and any $y\in[a,b]$, where $a<b$ and $(a,b)\in\mathbb{R}^{2}$ is in the support of $(Y^{L},Y^{U})$, there exists some $P^{\prime}\in\mathcal{P}_{I}$, such that $y\in L^{\prime}(\lambda)$, where $L^{\prime}(\lambda)$ is the upper level set for the density $p^{\prime}(y)$ of $P^{\prime}$.
Proposition 2.1 states that for any point $y\in[a,b]$ such that $(Y^{L},Y^{U})$ has positive density at $(a,b)\in\mathbb{R}^{2}$, we can find a random variable $Y^{\prime}$ that is observationally equivalent to $Y$ under censoring, and this $Y^{\prime}$ admits a level set that contains $y$ for any threshold $\lambda$. This is possible because one can adjust the density of $Y^{\prime}$ at any point $y$ satisfying the stated conditions to be arbitrarily large while ensuring that this density remains within the identified set $P_{I}$.
The following two lemmas characterise the identified set of the conditional distribution of $Y$ given $X$ under interval censoring. Lemma 2.2 relates the sharp identified set of the conditional distribution of $Y$ given $X$ to the conditional distribution of $(Y^{L},Y^{U})$, which can be directly identified and estimated from the data, and provides a feasible lower bound for the restriction in Equation (4). Lemma 2.3 shows that the lower bound is tight, and we are not unnecessarily conservative in constructing the oracle prediction set.
Lemma 2.2 (Theorem SIR-2.3 in Molinari, (2020)).
@@
Lemma 2.2 suggests that we could implement a feasible version of Equation (4), by replacing the optimisation constraint $\inf_{P\in\mathcal{P}_{I}}P\left(Y\in C(x)\mid X=x\right)\geq 1-\alpha$ with
$P\left([Y^{L},Y^{U}]\subset C(x)\mid X=x\right)\geq 1-\alpha.$
This implies the original constraint, since for $C\in\mathcal{C}$, we can write $C=\sqcup_{m=1}^{M}[a_{m},b_{m}]$, and
@@ 3 Consistency
We first consider the estimation of the oracle prediction interval $C^{*}_{I}=[\tau_{0}(x),\tau_{1}(x)]$, for $x\in\mathcal{X}$. This analysis highlights the conditions required to ensure the consistency of the prediction set under different schemes of interval censoring. Subsequently, we will extend our analysis to the estimation of oracle prediction sets $C^{*}\in\mathcal{C}$ consisting of multiple intervals. To measure the difference between two sets $A,B\subset\mathbb{R}$, we use the volume of their symmetric difference, denoted by $\mu(A\bigtriangleup B)$. The symmetric difference is defined as $A\bigtriangleup B=(A\setminus B)\cup(B\setminus A)$. This is commonly used in the literature on level set estimation and conformal inference. Another commonly used notion of set distance is the Hausdorff distance. In Appendix D, we discuss the relationship between these two metrics and why the Hausdorff distance is unsuitable in certain cases.
@@
$\min_{t_{0}<t_{1}}{t_{1}-t_{0}}\quad\text{s.t.}\quad P_{n}(t_{0},t_{1};x)\geq 1-\alpha.$ (7)
@@
where $K_{h}(x)=\frac{1}{h}K\left(\frac{x}{h}\right)$ for a choice of kernel smoothing function $K(\cdot)$ and bandwidth parameter $h$, see Tsybakov, (2009) and Hansen, (2008). In this case, $\mathcal{X}_{n}$ can be an expanding subset of $\mathcal{X}$ at a suitable rate to achieve uniform consistency as in Hansen, (2008), Theorem 8 (see discussion in Appendix B). Another candidate is the distributional regression estimator (see, e.g., Foresi and Peracchi,, 1995; Chernozhukov et al.,, 2013, 2021).
@@
Assumption 3.3 is a smoothness assumption on $P(t_{0},t_{1};x)$. In particular, we are assuming that by considering a slightly larger interval $[\tau_{0}-\varepsilon,\tau_{1}+\varepsilon]$ we can strictly increase our coverage probability. Assumption 3.3 also assumes an upper bound on the density, which can be violated if the conditional distribution of $(Y^{L},Y^{U})$ is discrete, which might happen under a fixed censoring scheme. This assumption is related to the $\gamma$-exponent condition in the level set literature, see Polonik, (1995) and Lei and Wasserman, (2014).
@@
$\min_{t_{0}<t_{1}}t_{1}-t_{0}\quad\text{s.t.}\quad P_{n}(t_{0},t_{1};x)\geq 1-\alpha-\psi_{n},$ (9)
@@
$P\left(\sup_{x\in\mathcal{X}_{n}}\sup_{t_{0}\leq t_{1}}\left\lvert P_{n}(t_{0},t_{1};x)-P(t_{0},t_{1};x)\right\rvert>\psi_{n}\right)=o(1).$ (10)
@@
Suppose $(Y_{i}^{L},Y_{i}^{U})=(0,1)$ or $(0,2)$ with equal probability without predictor information $X$. The oracle prediction set for a miscoverage level $\alpha=0.5$ (see Equation 4) is $C^{*}=[0,1]$. (Notice that $\inf_{P\in\mathcal{P}_{I}}P\left(Y\in C^{*}\right)=0.5$ and a shorter interval would not achieve the desired miscoverage level.) If we solve the exact empirical analogue of the oracle prediction set with the natural estimator $P_{n}(a,b)=n^{-1}\sum\mathbb{I}\left\{[Y_{i}^{L},Y_{i}^{U}]\subset[a,b]\right\}$ as in Equation 7, we obtain a prediction interval
$\check{C}=\left[0,1+\mathbb{I}\left\{\bar{Y}^{U}>1.5\right\}\right],$
where $\bar{Y}^{U}=\frac{1}{n}\sum_{i}Y_{i}^{U}$. Since the lower bound for $(Y_{L},Y_{U})$ is 0 with probability one, it is easy to see that the lower bracket of $\check{C}$ is 00. Because
$P_{n}(0,b)=\begin{cases}0&\quad\text{if}\quad 0<b<1\\ n^{-1}\sum\mathbb{I}\left\{Y_{i}^{U}=1\right\}&\quad\text{if}\quad 1\leq b<2\\ 1&\quad\text{if}\quad b\geq 2\end{cases},$
the upper bracket of $\check{C}$ is $2$ if $P_{n}(0,1)<0.5$, i.e. there are less observations of $Y_{i}^{U}=1$ than $Y_{i}^{U}=2$ in the sample, which is equivalent to having the average $\bar{Y}^{U}>1.5$. Since $\bar{Y}^{U}$ has a scaled Binomial distribution $\bar{Y}^{U}\sim 1+n^{-1}\text{Binom}(n,0.5)$, and the median for a $\text{Binom}(n,p)$ distribution is between $\lfloor np\rfloor$ and $\lceil np\rceil$, we have that $P(\bar{Y}^{U}>1.5)\to 0.5$ as $n\to\infty$. This means that $P(\check{C}=[0,2])\to 0.5$ as $n\to\infty$ and $\check{C}$ is not consistent. On the other hand, for any $0.50<\psi_{n}<$ 0.5, $P_{n}(0,1)<0.5-\psi_{n}$ is equivalent to $\bar{Y}^{U}>1.5+\psi_{n}$, which has probability approaching 00 by the central limit theorem if $\sqrt{n}\psi_{n}\to\infty$ as the sample size $n\to\infty$. Solving the modified optimisation problem in (9) will then lead to a consistent estimator of the oracle prediction set in this example. Also note that the condition in (10) is a generalisation of the condition $\sqrt{n}\psi_{n}\to\infty$.
@@
  1. 1.
    For any $\varepsilon>0$ and $P_{X}$-almost surely $x\in\mathcal{X}$, there exists $\delta=\delta(\varepsilon)>0$, such that for any set ${C}^{\prime}(x)\in\mathcal{C}$ with $\mu({C}^{\prime}(x))\leq\mu(C^{*}(x))$ and $\mu(C^{*}(x)\triangle C^{\prime}(x))>\varepsilon$, we have $P([Y^{L},Y^{U}]\subset{C}^{\prime}(x)\mid X=x)<1-\alpha-\delta$.
  2. 2.
    The oracle $C^{*}(x)\in\mathcal{C}_{\bar{M}}$, where $\mathcal{C}_{\bar{M}}=\{\sqcup_{m=1}^{M}[a_{m},b_{m}]:M^{\prime}\leq\bar{M}<\infty\}$, for some $\bar{M}<\infty$ and all $x\in\mathcal{X}$.
@@
4 Conformal inference with valued outcomes
@@
In particular, we will achieve finite-sample validity using the method of split conformal inference (see Romano et al., (2019) and Lei et al., (2018)). One feasible split conformal inference procedure for the oracle prediction interval $C^{*}_{I}$ in our setting can be described as follows. We first split the sample $\mathcal{D}$ into a training set $\mathcal{D}_{1}=\left\{Z_{i}:i\in\mathcal{I}_{1}\right\}$ and a calibration set $\mathcal{D}_{2}=\left\{Z_{i}:i\in\mathcal{I}_{2}\right\}$. We calculate the estimator $\hat{C}_{I}(x)=[\hat{\tau}_{0}(x),\hat{\tau}_{1}(x)]$ for the oracle prediction set with the training set for $x\in\mathcal{X}$ as in Equation 7. The following scores $s_{j}=s(Y^{L}_{j},Y^{U}_{j},X_{j};\hat{\tau}_{0},\hat{\tau}_{1})$ are then computed for each $j\in\mathcal{I}_{2}$ in the calibration set, given $\hat{C}_{I}(x)=[\hat{\tau}_{0}(x),\hat{\tau}_{1}(x)]$, where
@@
This score function is a modified version of the score function used in the case of conformalized quantile regression with no interval censoring studied in Romano et al., (2019). The score function is negative when the interval $[y^{L},y^{U}]$ is contained in $[t_{0}(x),t_{1}(x)]$ and positive otherwise. By allowing for negative scores, the conformal prediction set $\tilde{C}_{I}(x)$ penalises both overcoverage and undercoverage.
@@
where $q\left(\mathcal{S};\zeta\right)$ denotes the $\zeta$-quantile of the set $\mathcal{S}$ of real numbers. The conformal prediction set is then
@@
Intuitively, when our estimated prediction set $\hat{C}_{I}$ is wider than necessary, the intervals $\{(Y_{j}^{L},Y_{j}^{U}),j\in\mathcal{I}_{2}\}$ tend to be contained within $\hat{C}_{I}$, leading to a higher proportion of negative scores in the calibration set. As a result, the chosen quantile $\vartheta_{1-\alpha}$ is negative, and the conformal prediction interval adjusts for the over-coverage of $\hat{C}_{I}$. The correction for under-coverage follows the same logic.
@@
Notice that $s_{j}\leq 0$ if and only if $[Y^{L}_{j},Y^{U}_{j}]\subset[t_{0m},t_{1m}]$ for some $m\leq M$. We can then compute $\vartheta_{1-\alpha}=q(\{s_{j}:j\in\mathcal{I}_{2}\};(1-\alpha)(1+1/\left\lvert\mathcal{I}_{2}\right\rvert))$. The conformal prediction set is then given by  [^2] Here, the endpoints of the estimated $\hat{C}$ are adjusted uniformly using the same $\vartheta_{1-\alpha}$. A more sophisticated adjustment method is discussed in Appendix C, where we propose a minimal Euclidean distance adjustment. We conjecture that this approach could improve the convergence rate of the conformal prediction set. The optimal choice of conformity score remains an interesting question.
@@
Under the assumption that $\{(X_{i},Y_{i},Y_{i}^{L},Y_{i}^{U}):i\in\mathcal{I}\cup\{n+1\}\}$ are i.i.d. [^3] In order to perform conformal inference, it is sufficient that the sample satisfies the weaker condition of exchangeability. Here independence is assumed for simplicity and for establishing the asymptotics for the nonparametric estimators we will use. Barber et al., (2023) studies conformal inference beyond exchangeability and is applicable for time series data. Our results can be readily extended to the time series settings combined with their procedure. , the conformal prediction set $\tilde{C}$ in Equation 13 is finite-sample marginally valid, that is, for all $P$,
@@
Given a consistent estimator $\hat{C}$ for the oracle prediction set, if $s^{*}=s\left(Y^{L},Y^{U},X;C^{*}\right)$ has positive density in a neighbourhood around 00, then the conformal prediction set $\tilde{C}$ constructed as above, for an i.i.d. sample $\mathcal{D}=\left\{\left(X_{i},Y_{i}^{L},Y^{U}_{i}\right):i\in\mathcal{I}\right\}$ which is split into training $\mathcal{I}_{1}$ and calibration sets $\mathcal{I}_{2}$, we have
@@
However, it is impossible to construct a prediction set that is finite-sample conditionally valid and consistent for the oracle prediction set (Lei and Wasserman,, 2014; Barber et al.,, 2021; Gibbs et al.,, 2024), and hence we consider the weaker notion of local validity. Let $\mathcal{A}=\left\{A_{k}:k\leq K\right\}$ be a partition of $\mathcal{X}$, a prediction interval ${C}_{\text{loc}}(x)$ is finite-sample locally valid with respect to $\mathcal{A}$, if for all $P$,
$\operatorname{P}\left(Y_{n+1}\in{C}_{\text{loc}}(X_{n+1})\mid X_{n+1}\in\mathcal{A}_{k}\right)\geq 1-\alpha.$
@@
the local conformal prediction set is
$\tilde{C}_{\text{loc}}(x)=\bigcup_{m=1}^{M}\left[\hat{\tau}_{0m}(x)-\sum_{k\leq K}\mathbb{I}\left\{x\in A_{k}\right\}\vartheta_{1-\alpha,k},\hat{\tau}_{1m}(x)+\sum_{k\leq K}\mathbb{I}\left\{x\in A_{k}\right\}\vartheta_{1-\alpha,k}\right].$
The finite-sample local validity is a result of exchangeability, similar to Lei and Wasserman, (2014).
@@
Suppose $\{(X_{i},Y_{i}^{L},Y_{i}^{U}):i\in\mathcal{I}\}$ are i.i.d. Let $\mathcal{A}$ be a partition of $\mathcal{X}$. Then the local conformal prediction set $\tilde{C}_{\text{loc}}(x)$ is finite-sample locally valid with respect to $\mathcal{A}$.
In the next section, we demonstrate the performance of the conformal and local conformal prediction sets in a set of numerical experiments.